OpenStudy (darkbluechocobo):

A gasoline-driven lawnmower has a blade that extends out 1.5 ft from its center. The tip of the blade is traveling at the speed of sound, which is 1,100 feet/sec. What is the angular velocity of the tip of the blade? Express your answer in revolutions per minute (rpm). Use 3.14 as an approximation for π and round your answer to the nearest rpm.

OpenStudy (mathmath333):



OpenStudy (mathmath333):

what do the think the circumference for the blade would be ?

OpenStudy (darkbluechocobo):

2(3.14)1.5= 9.4

OpenStudy (darkbluechocobo):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{circumference =2\pi r \hspace{.33em}\\~\\ C=2\pi 1.5 \hspace{.33em}\\~\\ C=9.42~~feet \hspace{.33em}\\~\\ \normalsize \text{one revolution = circumference one time }\hspace{.33em}\\~\\ 1~revolution=9.42~~feet \hspace{.33em}\\~\\ 1~feet=\dfrac{1}{9.42}~~revolution \hspace{.33em}\\~\\ angular ~~velocity=1,100 ~~feet/sec \hspace{.33em}\\~\\ =1,100\times \dfrac{1}{9.42}~~revolution /sec \hspace{.33em}\\~\\ =1,100\times \dfrac{1}{9.42}\times 60~~revolution /min \hspace{.33em}\\~\\ =7006~~revolution /min \hspace{.33em}\\~\\ }\end{align}\)